I am stuck on finding a quadratic model for (1,-2), (2,-2), (3,-4). I do not know how to put x numbers and y numbers in the formulas and how to solve the answer?
Answer by stanbon(75887)

I am stuck on finding a quadratic model for (1,-2), (2,-2), (3,-4). I do not know how to put x numbers and y numbers in the formulas and how to solve the answer?
--
Form: ax^2 + bx +c = y
Procedure: Substitute each x/y pair into the form to give you
three equations with three unknown variables: a, b, and c.
Solve the system of three equation for a, b, and c.
---------------------------
(1,-2) gives a + b + c = -2
(2,-2) gives 4a+2b + c = -2
(3,-4) gives 9a+3b + c = -4
-----------------------------------
Solve the system by any mehod you know to get:
a = -1
b = 3
c = -4
-------------
Equation:
y = -x^2 + 3x -4
=======================
Cheers,
Stan H.